# Tuning Colloidal Reactions

### v1

@@ Abstract
The precise control of complex reactions is critical for biological processes ranging from cell division to metabolism. Synthetic analogues of living materials suffer from our inability to tune chemical reactions with precise outcomes. Here, we leverage differentiable simulators to design nontrivial reaction pathways in colloidal assemblies. By optimizing interactions between reactants and substrates, we achieve controlled disassembly of octahedral and icosahedral shells. As a potential engineering target, we design a reaction that provokes the release of a small particle trapped in a shell.
Both living and non-living physical systems exhibit complex, dynamical behavior, ranging from repair, locomotion, and catalysis. Fundamentally, these complex dynamical behaviors arise from sequences of reactions, in which a set of substances (i.e. the reactants) are transformed into a set of different substances (i.e. the products).
A rich body of theoretical work aims to characterize and tune systems of interacting agents spanning a range of system descriptions [1, 2, 3]. There have been exciting advances in creating tunable reactions that are experimentally realizable [4, 5, 6, 7, 8, 9]. However, the models that provide an experimentally-relevant level of detail are typically bespoke and application-specific, necessitating entirely new research programs for each new reaction. Conversely, general theoretical models have provided deep insights into the basis of self-limiting growth [10, 11], self-replication [12],catalysis [13, 14], and much more [15, 16, 17, 18, 19]. These models, however, are largely too abstract to inform experimental design. Moreover, many reactions that yield desirable dynamical behavior are incredibly finely-tuned, making engineering these reactions untenable with most traditional design approaches. To design precise reactions, it is necessary to carry out inverse design, whereby one designs components and their interactions to achieve a given reaction. While inverse design has successfully applied to self assembly [20, 21, 22, 23, 24], its application to reaction pathways is much more difficult owing to the need to choose design parameters to favor particular dynamical trajectories. The advent of differentiable simulators [24], powered by software libraries developed for machine learning [25], has opened up the possibilty of directly designing reactions.
@@
Here, we use differentiable molecular dynamics (MD) to carry out the inverse design of complex reactions. The differentiability makes it possible to compute analytical gradients with respect to trajectories during the interaction of components and perform gradient-based optimization with respect to parameters to achieve a target reaction. To demonstrate the ability framework to design nontrivial reactions, we consider the controlled disassembly of colloidal structures, whereby one particle is extracated from an otherwise complete shell of colloidal particles. Disassembly is central to the dynamic functions of living systems, such as defect repair, self-replication, and catalysis. Most existing examples of controlled disassembly in synthetic systems rely on external forcing to drive the disassembly process [26, 27, 28, 29], which provides a direct pathway to tuning behavior. However, for many engineering applications including those inherent to living systems, the use of external fields is limiting. Alternatively, living systems typically rely on local energy consumption (e.g. biological enzymes consuming ATP) rather than global fields, but the synthetic design of these systems is significantly more complex.
@@
In this paper, we design for the controlled disassembly of both an octahedral shell and an icosahedral shell. Importantly, our disassembly mechanism does not rely on external forcing and occurs without consuming energy. As a model for potential engineering applications, we apply our disassembly mechanism to release a small particle trapped inside an icosahedral shell. Controlled disassembly serves as a striking example of a complex reaction because the reaction requires a finely-tuned interaction energy to keep the remaining shell stabilized while performing the desired disassembly. Moreover, we consider a fixed substrate and only parameterize an external structure that acts on this fixed substrate, enabling control over disassembly without modifying critical components of the target reaction.
@@
Results.—We implement controlled disassembly example in a patchy particle system. Patchy particles have long been used to emulate interactions in soft materials[30, 31], and offer tremendous flexibility of designed interactions, especially when patches can be programmed to interact with each other with designed strength and range. Optimizing patchy particle systems to achieve specific behaviors has been made possible by the recent development of patchy particle simulations within a differentiable library [24].The ability to carry out this type of optimization within a molecular dynamics simulations opens up enormous possibilities for the design of novel reactions. In particular, we aim to remove a single particle from a shell composed of patchy particles without disrupting the remaining shell structure. We tune disassembly without changing any properties of the shell itself, by introducing an external reactant that interacts with the shell to disassemble it in the desired manner. We term the external reactant a “spider” due to the structure of the object.
@@
The shell is modelled as a collection of patchy particles where each patch corresponds to a contact with a neighboring particle. Patches interact via a Morse potential (, ) and vertex centers interact via soft-sphere repulsion (). Importantly, the geometry and interaction energy of a shell vertex are fixed throughout an optimization: only features of the spider are optimized.
@@
The spider is modeled as a rigid body composed of a ring of “base” particles and a “head” particle that sits above the ring along its symmetry axis. The base particles are connected by repulsive bars and the head is connected to the base particles particles by the same repulsive bars, making the entire structure a cage-like object that is open on one end. The head particle interacts with shell vertices via a Morse potential whereas base particles and connecting bars interact with shell vertices via soft sphere repulsion. Unlike the shell, the geometry and interaction energy of the spider are parameters of the optimization. See Figure 1C for an overview of this parameterization. All the interaction energies in our system are parameterized with simple, physics-based potentials.
We optimize over 8 parameters that characterize the geometry of the spider and its interaction with the shell via differentiable MD (see Figure 1D). Our loss term is constructed via two competing terms: one that rewards a final state in which the target vertex is extracted, and one that penalizes a strong interaction between the spider and non-target vertices. Consider a shell comprised of a collection of vertices where and for the octahedron and icosahedron, respectively. We seek to extract a target vertex from the shell while leaving the remaining shell intact. We can measure the degree to which is successfully extracted via the following expression:
@@
where denotes the Euclidean distance between vertices and . Note that the negative sign as we formulate our optimization problem to minimize the loss. Secondly, we can minimize the interaction energy between the spider and the non-target vertices of the shell,
@@
where represents the interaction energy between the spider head, , and a shell vertex as depicted in Figure 1C. For the total loss, we evaluate the interaction between the spider and the non-target shell vertices with respect to the initial state while we calculate the spider’s extraction of the target vertex with respect to the final state. We evaluate the remaining energy term on the initial state because this term is only well defined when evaluated on the same configuration at each iteration. In all simulations, the spider is initialized bound to the target vertex and the system is integrated for 1000 timesteps (see Appendix).
Octahedron.—We first demonstrate our method on an octahedron, a simple but nontrivial platonic solid. We explore two limits of our optimization procedure: weak and strong initial shell-spider interactions. Figure 2 summarizes our results.
@@
First, we perform an optimization where the spider is initialized to interact weakly with the shell vertices (, ). Under this weak initial interaction, the spider simply diffuses away from the shell at long timescales without extracting the target vertex. Then, throughout the course of the optimization, we observe variable changes consistent with increasing the interaction between the spider head and the shell – increases, the head height decreases, and the head particle radius increases. However, we also observe less intuitive changes – increases (decreasing the range of the Morse potential), the shell base radius decreases, and the base particle radius increases. This suggests tightly coupled, nontrivial parameter changes drive the extraction of the target vertex while maintaining minimal interaction with the remaining shell.
@@
Next, we performed an optimization in the opposite limit in which the spider is initialized to interact strongly with the shell (, ). Initially, this interaction is so strong that the spider not only extracts the target vertex but it also disrupts the remaining shell. This disruption can be seen in the large value of the remaining energy loss term, which penalizes the energy between the spider head and non-target vertices. Throughout the optimization, we observe variable changes consistent with tuning the interaction strength to maintain extraction while minimizing off-target interactions – decreases, increases, the head radius decreases, the head height increases, and the base particle radius increases. When evaluated on longer simulations, the converged parameter set also achieves diffusion of the spider-vertex complex from the remaining, undisturbed shell.
While only single optimizations are presented in Figures 2 and 3, we additionally performed ensembles of optimizations over different initial seeds for the random number generator and over different random perturbations to the initial parameter in both the high and low energy regimes. Despite the wide variance in initial parameters, we observe low variance in the converged parameter values in all cases (see Figure 4). This may suggest that the solution set is narrow.
@@
Contrasting these two optimization regimes reveals the inherent delicacy in tuning the spider to achieve extraction of a target vertex and subsequent diffusion of the vertex away from the shell. The spider-shell interaction must be sufficiently strong to extract the target vertex, but be simultaneously weak enough to not disturb the non-target vertices and to diffuse away from the shell within the timescale of our simulations. This tension can be seen in the behavior of the loss terms in each optimization. In the weak-interaction limit, the term penalizing interactions with non-target vertices remains negligible while the extraction term drives optimization; in the strong-interaction limit it is the same energy-penalizing term that dominates the loss.
Icosahedron and Controlled Release.—We now increase in complexity to a model of a biologically relevant function: controlled release of a particle from an icosahedral shell. We initialize our system with a small particle trapped inside a fully formed icosahedral shell. We then optimize a similar spider structure to perform controlled disassembly of the icosahedral shell, resulting in the controlled release of the small particle. This serves as a toy example of a potential target for engineering applications, such as drug delivery via a viral shell.
To match the symmetry of the icosahedron, we use a spider with five legs rather than four. Otherwise, optimization proceeds in the same fashion as for the octahedral shell. We consider the same two limits: one with high interaction strength between the spider and the shell and one with low interaction strength between the spider and the shell (see Figure 3). The same delicate tradeoff between extraction and diffusion manifests with the icosahedron, and our automatic differentiation-based optimization deftly navigates this tradeoff.
@@
Discussion.—In this Letter, we have demonstrated the design of nontrivial reactions via differentiable simulation. We considered the case of controlled disassembly, in which there is an inherent tension between initiating disassembly and maintaining the integrity of the remaining substructure. We first consider the controlled disassembly of an octahedral shell composed of patchy particles and show how the parameters governing an external structure (i.e. the “spider”) can be finely tuned to minimize a loss function representing this inherent tension. We then increase in complexity, demonstrating the controlled release of a small particle from an icosahedral shell.
@@
By optimizing directly with respect to the numerically integrated dynamics, our method is general enough to study a wide range of systems. Foremost, our method may enable the experimental realization of theoretical models that were otherwise limited by an inability to finely tune interaction energies. For example, Ref. [12] introduces a model of self-replicating colloidal clusters in which kinetic traps can be avoided by tuning the interaction energies, but dissociation of a new cluster from its parent (a necessary step for replication) required an artificial trigger event in numerical simulations. In contrast, our designed parameters lead to spontaneous dissociation of the spider-vertex complex away from the remaining shell. The computational flexibility of the method could also easily allow restricting the parameter regime to experimentally realizable interactions. This could be done for DNA coated colloids, for instance, by optimizing the DNA sequences that define the interaction strength [32, 33].
@@
A number of numerical instabilities can arise while optimizing for the types of reactions we consider. The primary limitation we observe is that gradients become unstable and very large for long simulations. There are several possible approaches to reducing instability in gradients of long simulations. One standard method to mitigate such instabilities in the context of differentiable programming is gradient clipping [34, 35]. One could also decrease the total number of timesteps by training an emulator to resolve the dynamics with a larger timestep than is otherwise possible with standard integrators, following similar work for deterministic systems [36, 37, 38]. An alternative approach to decreasing the number of timesteps would be to integrate differentiable simulations with enhanced sampling methods. As a result, one could sample low probability events without the need for long simulation times.
@@
We rely on gradient-based optimization due to its scalability and performance. The method scales naturally to larger systems since the calculation of a gradient via automatic differentiation only requires a single simulation, while at the same time reverse-mode automatic differentiation scales efficiently with the number of parameters [39]. Calculating the analytical gradient with respect to a simulation explicitly captures interdependency between parameters, essential to efficiently tuning complex behavior.
The experimental realization of theoretical models of complex dynamical behavior has long been limited by the ability to finely tune the parameterization of such models. This method serves as a potential bridge between these two worlds, permitting the delicate tradeoffs inherent in parameterizations of complex reactions to be navigated efficiently using physics-based potentials and dynamics.

### v2

@@ Abstract
The precise control of complex reactions is critical for biological processes yet our inability to design for specific outcomes limits the development of synthetic analogues. Here, we leverage differentiable simulators to design nontrivial reaction pathways in colloidal assemblies. By optimizing over external structures, we achieve controlled disassembly and particle release from colloidal shells. Lastly, we characterize the role of configurational entropy in the structure via both forward calculations and optimization, inspiring new parameterizations of designed colloidal reactions.
Both living and non-living physical systems exhibit complex dynamical behavior, ranging from repair to locomotion to catalysis. Fundamentally, such behaviors arise from sequences of reactions, in which a set of substances (i.e. the reactants) are transformed into a set of different substances (i.e. the products). A rich body of work aims to characterize and tune systems of interacting agents spanning a range of system descriptions, both theoretically [1, 2, 3] and experimentally [4, 5, 6, 7, 8, 9]. However, for many critical processes in biological systems (e.g. DNA synthesis, protein folding), the components themselves cannot be changed. Instead, to modify these processes, researchers often introduce an external structure (e.g. competitive inhibitors for enzyme inhibition, protein folding chaperones). Thus far, the design of such structures has been bespoke and application-specific, necessitating entirely new research programs for each new reaction. For example, while some general theoretical models have provided deep insights into catalysis [10, 11], they are largely too abstract to inform experimental design.
To overcome current limitations and tune reactions through the design of external agents, we carry out inverse design whereby we optimize the geometry and interactions of such components to achieve a target reaction. While inverse design has been successfully applied to self-assembly [12, 13, 14, 15, 16], inverse-designing reaction pathways remains a challenge because design parameters must be chosen to favor particular dynamical trajectories. The advent of differentiable simulators [16], powered by software libraries developed for machine learning [17], has opened up the possibility of directly designing reactions as the gradient of numerical procedures with respect to control parameters can be computed efficiently.
@@
Here, we design complex reactions using differentiable molecular dynamics (MD) and gradient-based optimization. As an example of a nontrivial reaction, we consider the controlled disassembly of colloidal structures, whereby a particle is extracted from an otherwise complete shell of colloidal particles. Disassembly is central to the dynamic functions of living systems, such as defect repair, self-replication, and catalysis. Existing examples of controlled disassembly in synthetic systems often rely on external forcing to drive the disassembly process [18, 19, 20, 21], which provides a direct pathway to tuning behavior. However, for many engineering applications including those inherent to living systems, the use of external fields is limiting. On the other hand, controlled disassembly in living systems typically relies on local energy consumption (e.g. biological enzymes consuming ATP) rather than global fields, but the synthetic design of these systems is significantly more complex.
@@
Inspired by the symmetry of many viral capsids [22, 23], we design for the controlled disassembly of icosahedral shells. We consider a fixed shell and only parameterize an external structure that acts upon it, enabling control over disassembly without modifying critical components of the reaction. Importantly, our disassembly mechanism is entirely passive and does not rely on external forcing. As a model for potential engineering applications, we apply our mechanism to provoke the release of a target small particle initially trapped inside the shell. Controlled disassembly serves here as a striking example of a complex reaction because the reaction requires a finely-tuned interaction energy to keep the remaining shell stabilized while still performing the desired particle extraction. We start from a rigid structure and thereafter proceed to quantify the role of flexibility in the structure by computing free energy landscapes both for predefined extrema and for structures optimized via a chosen parameterization of configurational entropy, opening the door to novel designed reactions.
@@
Results.—We implement controlled disassembly in a colloidal patchy particle system. Patchy particles have long been used to emulate interactions in soft materials [24, 25] and offer tremendous tunability in designed interactions. Optimizing said systems to achieve specific behaviors has been made possible by the recent development of patchy particle simulations within a differentiable library [16]. In particular, we aim to remove a single particle from a shell composed of patchy particles in a controlled manner without disrupting the remaining shell structure. To that end, we tune disassembly without changing any properties of the shell itself. Instead, we introduce an external structure that interacts with the shell to disassemble it in the desired manner. We term the external structure a “spider” due to its geometry.
@@
The shell is modelled as a collection of patchy particles forming an icosahedron where each patch corresponds to a contact with a neighboring particle. Each patchy particle consists of a central sphere and a set of rigidly attached patches. Patches interact via a Morse potential (, ) and particles interact via soft-sphere repulsion (). Importantly, the geometry and interaction energy of shell-comprising particles are fixed throughout the optimization. Although we focus on the disassembly of the icosahedron, our framework can be easily adapted for octahedral shells (see Appendix).
@@
For the spider, we consider several different models with a set of core similarities. All models contain a ring of “base” particles and a “head” particle that sits above the ring along its symmetry axis. The head is connected to the base particles by repulsive bars, making the entire structure a cage-like object that is open on one end. An attractive particle type (either the head or a third particle species) interacts with the shell-comprising particles via a Morse potential whereas base particles and connecting bars interact with shell particles via soft sphere repulsion. Unlike the shell, the geometry and interaction energy of the spider are parameters of the optimization. See Figure 1C for an overview of this parameterization. All interaction energies in our system are parameterized with simple, physics-based potentials.
Given a specified parameterization for the spider geometry and interactions, we run an ensemble of differentiable molecular dynamics simulations (see Figure 1D and Appendix). To focus our optimization procedure on the challenges specific to disassembly, we initialize the spider bound to the shell and therefore ignore the period in which the spider is freely diffusing. We optimize over the parameters that characterize the geometry of the spider and its interaction with the shell (8 parameters for the optimizations in Figure 2). To optimize our system, we perform gradient descent to minimize a loss function. The loss is constructed from two competing terms: one that rewards a final state in which the target particle is extracted, and one that penalizes a strong interaction between the spider and non-target particles. The second term, which we refer to as the “remaining energy” term, tends to reward pathways in which the remaining shell holds its shape after the spider extracts the target particle.
We formalize the loss function as follows. Consider an icosahedral shell comprised of a collection of particles where . We seek to extract a target particle from the shell while leaving the remaining shell intact. We can measure the degree to which is successfully extracted via the following expression:
@@
where denotes the Euclidean distance between particles and . Note the negative sign as we formulate our optimization problem to minimize the loss. Next, we minimize the interaction energy between attractive site(s) and non-target shell particles:
@@
where denotes the set of attractive sites and represents the interaction energy between the attractive site and a shell particle . For the spider depicted in Figure 1, the head particle is the only attractive site. We calculate the ‘remaining energy’ term, , of the total loss with respect to the initial configuration, i.e. the first timestep of the simulation. All other terms depend on the dynamics of the system, so we evaluate them on the final state. In all simulations, the spider is initially bound to the target particle and we integrate the system for 1000 timesteps (see Appendix).
Rigid Spider.– We begin with a minimal conception of the spider: a rigid body consisting of only a head particle and five base particles which reflect the five-fold symmetry of the icosahedron. The base particles are attached to each other by rigid bars, forming a cage-like structure that is open on one end. The attractive interactions between the spider and the icosahedron are restricted to interactions between the head particle in the spider and the patches on the icosahedral vertices.
@@
We explore two limits of our optimization procedure (Figure 2). First, we perform an optimization where the spider is initialized to interact weakly with the shell particles (, ). In this limit, the spider simply diffuses away from the shell at long timescales without extracting the target particle. Initially, we observe variable changes consistent with increasing the interaction between spider head and shell to achieve particle extraction: increases, the head height decreases, and the head particle radius increases. In the following iterations, we observe parameter changes focused on maintaining extraction while reducing the interaction strength between the spider and the rest of the shell. The head height increases, consistent with minimizing the remaining energy, but to maintain particle extraction, decreases (increasing the range of the Morse potential). This suggests that tightly coupled, nontrivial parameter changes drive extraction while maintaining minimal interaction with the remaining shell.
@@
Next, we perform an optimization in the opposite limit in which the spider is initialized to interact strongly with the shell (, ). Initially, this interaction is so strong that the spider not only extracts the target particle but it also disrupts the remaining shell. This can be seen in the large value of the remaining energy loss term, which penalizes the energy between the spider head and non-target particles. Throughout the optimization, we observe variable changes consistent with tuning the interaction strength to maintain extraction while minimizing off-target interactions: decreases, increases, the head radius decreases, the head height increases, and the base particle radius increases. When evaluated on longer simulations, the converged parameter set also achieves spontaneous diffusion of the spider-particle complex from the remaining undisturbed shell. Note that neither changes in the random seed nor perturbations to the initial parameters consistently yield similar optimized parameters (see Appendix).
@@
Contrasting the high and low energy optimization regimes reveals the inherent delicacy in tuning the spider to achieve extraction and subsequent diffusion from the shell. The spider-shell interaction must be sufficiently strong to extract the target particle, but simultaneously weak enough to not disturb non-target particles and to diffuse away from the shell within the timescale of our simulations. This tension is reflected in the behavior of the loss terms in each optimization. Overall, in the weak-interaction limit, the term penalizing interactions with non-target particles remains negligible while the extraction term drives optimization; in the strong-interaction limit it is the same energy-penalizing term that dominates the loss. Our optimized reactions represent a notion of balance that is necessary for biologically relevant functions, e.g. the controlled release of a particle from a closed shell (see simulation in Figure 2C). This serves as a toy example of a potential target for engineering applications, such as drug delivery via a viral shell.
Flexible Spider.– The configurational entropy of the spider can serve as a control knob for tuning reactions. While the optimized results in Figure 2 highlight that spider geometry dramatically impacts its performance, the rigid formulation does not access configurational entropy. Here, we define a modified form of our spider permitting versions with varying degrees of flexibility. Rather than the head serving as the sole attractive site, we introduce a ring of attractive sites consisting of one site per spider leg positioned between the base and head particles (Figure 3A). In this way, bonds at the spider base connecting individual legs can be made flexible or removed entirely. The resulting fluctuations due to leg flexibility directly change the interaction strength between spider and extracted particle and thus modulate the entropic contribution to the interaction. In this scheme, the head particle only interacts repulsively with the icosahedron.
Instead of considering the probability of extraction, we focus on the release of an already extracted particle since this process is likely to be strongly influenced by configurational entropy. We reason that increased entropy in the extracted state (i.e. the extracted particle bound to the spider) would favor particle release compared to the fully rigid spider because fluctuations in the spider configuration would reduce the effective attraction. To test this hypothesis, we quantitatively measure free energy differences corresponding to particle release between three versions of the modified spider with varying flexibility: (i) a fully rigid spider with fixed bonds between all adjacent base particles, (ii) a partially flexible spider resulting from the removal of two base bonds, and (iii) a fully flexible spider via the removal of all base bonds.
We compute the free energy of release using each of the three models. The distance from the extracted particle to the head directly relates to its release. We therefore use this metric as the order parameter for free energy calculations (Figure 3B). We compute the free energy diagrams for each spider using the Weighted Histogram Analysis Method (WHAM) and use a fixed set of parameters for the spider geometry and interaction determined via a single optimization with the fully rigid spider (see Appendix). As expected, the more flexible the geometry, the more favorable the released state: there is a smaller change in free energy between the attached and released states for more flexible geometries (Figure 3C).
Next, we optimize configurational entropy directly. We define a spider in which all pairs of (i) adjacent and (ii) next-nearest neighbor base particles are connected with springs whose spring constants are free parameters (Figure 3D). To bias the optimization procedure towards a spider with an increased likelihood of release, we define an additional loss term representing the total interaction energy between attractive sites and extracted particle:
The weaker the interaction, the easier the release. Here, we optimize over a longer (1500 step) simulation than in the rigid case. We average the new loss term over states sampled from the final 500 simulation steps, while maintaining extraction within the first 1000 steps (see Appendix). To give the optimization algorithm more freedom to promote release, we rescale the loss describing extraction such that it changes minimally beyond a specified maximum value. The optimization algorithm can then reduce extraction efficiency without penalty.
We optimize these spring constants using the modified loss function while keeping all other parameters fixed. We initialize all spring constants to be the same value (i.e. ) and fix the spider geometry and interaction parameters to those used in Figure 3C. The optimized solution has a wider well but maintains the same well depth. As a result, the free energy difference between the extracted and released states is lower in the optimized configuration than in the initial one at intermediate distances. Interestingly, the optimization procedure naturally converges to a solution with an asymmetric distribution of spring constants. Directly tuning this asymmetry is a promising avenue for future work.
@@
Discussion.— In this Letter, we achieve nontrivial reactions via designed external structures. We consider the case of controlled disassembly of an icosahedral shell composed of patchy particles, in which there is an inherent tension between initiating disassembly and maintaining the integrity of the remaining substructure. We show how the parameters governing a rigid external structure can be finely tuned to minimize a loss function representing this tension. We find that the optimized spider provokes particle release. We then add configurational entropy by introducing a flexible spider geometry, and quantify the influence of flexibility by comparing free energy landscapes for varying degrees of flexibility. Our framework naturally accommodates parameterizations of configurational entropy. Upon optimization, a spider with asymmetrically flexible base legs favors release over the initialized uniform configuration.
@@
Since we optimize directly with respect to the numerically integrated dynamics, our method is general enough to study a wide range of systems. Foremost, it may enable experimental realizations of theoretical models that were otherwise limited by an inability to finely tune interaction energies. For example, Ref. [26] introduces a model of self-replicating colloidal clusters in which kinetic traps can be avoided by tuning the interaction energies, but dissociation of a new cluster from its parent (a necessary step for replication) required an artificial trigger event in numerical simulations. In contrast, our designed parameters lead to spontaneous dissociation of the spider-particle complex away from the remaining shell. The computational flexibility of the method could also easily enable users to restrict the parameter regime to experimentally realizable interactions. This could be done for DNA coated colloids, e.g., by optimizing the DNA sequences that define the interaction strength [27, 28, 29].
@@
While optimizing for such types of reactions, numerical instabilities may arise. The primary limitation we observe is that gradients become unstable and very large for long simulations. There are several possible approaches to reducing instability in gradients in such cases. One standard method to mitigate instabilities in the context of differentiable programming is gradient clipping [30, 31]. One could also decrease the total number of timesteps by training an emulator to resolve the dynamics with a larger timestep than is otherwise possible with standard integrators, following similar work for deterministic systems [32, 33, 34]. An alternative approach would be to integrate differentiable simulations with enhanced sampling methods to sample low probability events without the need for long simulation times.
@@
We rely on gradient-based optimization due to its scalability and performance. Naturally, our method scales to larger and more complex systems since (i) gradient calculation via automatic differentiation only requires a single simulation, (ii) reverse-mode scales efficiently with the number of parameters [35], and (iii) the gradient explicitly captures interdependencies which is essential to efficiently tuning complex behavior. We envision that our approach and proposed design rules will be applicable to physical reactions beyond the colloidal regime.